a carpenter has 2 peices of wallpaper border to use in a house. one piece is 5 feet longer than the other. together the two pieces measure 105 feet. what is x, the length of the shorter piece of border?
Answer by queenofit(120)

Ok, first you know you have 2 pieces of boarder.
let x= the shorter piece
let x+5= the longer piece (only logical because it is 5ft longer)
set up your equation like this.
x+x+5=105
2x+5=105 (you can add the x's together)
2x+5-5=105-5 (you must get like terms together)
2x=100 (divide both sides by 2)
x=50
x+5=55
the shorter piece is 50ft and the longer piece is x+5 so it is 55ft.
hope this helps